calculate the kinetic energy per molecule of a gas at 0 degree celsius ?

Grade:11

2 Answers

Arun
4 years ago
Actually the average kinetic energy can easily be determined by the formula :
3/2 KT . Here K is the boltzmann constant which is equal to 1.38x10^-23 J/K
and T representing the temperature in kelvin 
0 degree celsius = 273 kelvin so by just putting the values in the formula
 
E = 3/2 * 1.38 * 10^-23 * 273
 
 
 
Vikas TU
4 years ago
Dear student 
Kinetic Molecular Theory formula is:
KE = 3/2 •kB•T
Where:
KE is the average kinetic energy of the gaseous molecule.
kB is the Boltzmann Constant (1.3806 x 10-23 m2 kg s-2 K-1)
T is the Temperature in Kelvin
This calculates the average kinetic energy of a gaseous molecule at a specific temperature.
E = 3/2 * 1.38 * 10^-23 * 273